# suppSupplementary References

### v1

@@ Abstract
Abstract: Raw nanopore signal analysis is a common approach in genomics to provide fast and resource-efficient analysis without translating the signals to bases (i.e., without basecalling). However, existing solutions cannot interpret raw signals directly if a reference genome is unknown due to a lack of accurate mechanisms to handle increased noise in pairwise raw signal comparison. Our goal is to enable the direct analysis of raw signals without a reference genome. To this end, we propose Rawsamble, the first mechanism that can 1) identify regions of similarity between all raw signal pairs, known as all-vs-all overlapping, using a hash-based search mechanism and 2) use these to construct genomes from scratch, called de novo assembly.
Our extensive evaluations across multiple genomes of varying sizes show that Rawsamble provides a significant speedup (on average by 16.36×16.36$\times$16.36 × and up to 41.59×41.59$\times$41.59 ×) and reduces peak memory usage (on average by 11.73×11.73$\times$11.73 × and up to by 41.99×41.$9$) compared to a conventional genome assembly pipeline using the state-of-the-art tools for basecalling (Dorado’s fastest mode) and overlapping (minimap2) on a CPU. We find that $36.57\%$ of overlapping pairs generated by Rawsamble are identical to those generated by minimap2. Using the overlaps from Rawsamble, we construct the first de novo assemblies directly from raw signals without basecalling. We show that we can construct contiguous assembly segments (unitigs) up to 2.7 million bases in length (half the genome length of E. coli). We identify previously unexplored directions that can be enabled by finding overlaps and constructing de novo assemblies.
@@ 1 Introduction
Nanopore sequencing technology [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23] can sequence long nucleic acid molecules of up to a few million bases at high throughput [24, 25, 26]. As a molecule moves through a tiny pore, called a nanopore, ionic current measurements, called raw signals, are generated [5]. Nanopore sequencing provides three unique key benefits.
@@
Existing works that analyze raw nanopore signals [30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 27, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66] mainly utilize deep learning mechanisms [30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47] to translate these signals into nucleotides, a process called basecalling. Basecalling mechanisms usually 1) are designed to use large chunks of raw signal data for accurate analysis [53] and 2) have high computational requirements [60, 26]. This can impose limitations to enable 1) accurate real-time analysis [53] and 2) portable sequencing with constrained resources [60].
To fully utilize the unique benefits of nanopore sequencing, it is necessary to analyze raw signals with 1) high accuracy and low latency for adaptive sampling and 2) low resource usage for portability and efficiency. To achieve this, several mechanisms focus on analyzing raw nanopore signals without basecalling [52, 27, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66]. [^1] We use the raw signal analysis term specifically for these mechanisms in the remainder of the paper. Although raw nanopore signal mapping to a reference genome is widely studied to achieve relatively accurate and fast mapping of raw signals [27, 53, 54, 57, 58, 59, 60, 62, 63, 64, 65, 66], none of these works without a reference genome.
When a reference genome is not available, a genome can be constructed from scratch, called de novo assembly construction[67]. To construct a de novo assembly, similar regions between all read pairs are identified, called all-vs-all overlapping, instead of identifying similarities between reads and a reference genome. [68, 26, 69]. However, an all-vs-all overlapping between raw nanopore signals remains unsolved due to several unique challenges [5, 70, 71, 72].
First, it is challenging to identify similarities between a pair of noisy raw signals as compared to a noisy raw signal and an accurate signal generated from a reference genome, an approach commonly used in earlier works for read mapping [54, 53, 58, 59, 57]. This is because converting reference genomes to their expected raw signal values is free from certain types of noise that raw nanopore signals contain (e.g., stochastic signal fluctuations [5] and variable speed of DNA molecules moving through nanopores [70, 71]), while none of the raw signals are free from such noise in all-vs-all overlapping.
@@
Our goal is to enable 1) raw signal analysis without a reference genome and 2) new use cases with raw nanopore signal analysis, such as assembly from overlapping raw signals, by addressing the challenges of accurate and fast all-vs-all overlapping of raw nanopore signals. To this end, we propose Rawsamble, the first mechanism that enables fast and accurate overlap finding between raw nanopore signals. The key idea in Rawsamble is to re-design the existing state-of-the-art hash-based seeding mechanism [58, 59] for raw signals with more effective noise reduction techniques and useful outputting strategies to find all overlapping pairs accurately, which we explain in three key steps.
@@
  • We propose the first mechanism that can find all-vs-all overlapping of raw signals without basecalling to enable new use cases for raw signal analysis, such as assembling raw signals.
  • We show that identifying overlaps with Rawsamble 1) is faster (on average 16.36×16.36$\times$16.36 × and up to 41.59×41.59$\times$41.59 ×) and 2) reduces peak memory requirements (on average 11.73×11.73$\times$11.73 × and up to 41.99×41.$9$) compared to the computational pipeline that includes the state-of-the-art basecalling running on a CPU with its fastest model followed by minimap2 to find overlaps. We evaluate the trade-offs between speed, accuracy, and hardware resources by running the basecaller on a CPU and GPU with fast and high-accuracy models.
  • We show that $36.57\%$ of overlapping pairs that Rawsamble generates are identical to the overlapping pairs generated by minimap2.
  • We report the first de novo assemblies ever constructed directly from raw signal overlaps without basecalling. We show that we can construct long unitigs up to 2.7 million bases in length for E. coli, which constitutes half the length of the genome.
  • We discuss new use cases that can be enabled with overlapping raw signals and constructing assemblies from raw signals, such as new directions for improving the accuracy and performance of basecalling.
@@ 2.1 Overview
Rawsamble is a mechanism to find overlapping pairs between raw nanopore signals (i.e., all-vs-all overlapping) for constructing de novo assemblies without basecalling, as shown in Figure 1. To achieve this, Rawsamble builds on the state-of-the-art raw signal mapper, RawHash2 [58, 59]. We provide the overview of RawHash2 in Supplementary Section A. Rawsamble extends RawHash2 to support all-vs-all raw signal overlapping in four key steps: First, to enable efficient and accurate indexing from the noisy raw signals ( 1), Rawsamble aggressively filters the raw input after the signal-to-event conversion to avoid nanopore-related errors. Second, to improve the accuracy of overlapping ( 2), Rawsamble adjusts the minimum chaining score to avoid false chains, ensuring that only high-confidence overlaps are considered. Third, to enable finding useful and long connections from many overlaps, Rawsamble adjusts the output strategy such that 1) all chains rather than only the best chain are reported and 2) cyclic overlaps are avoided. Fourth, we introduce a new assembler to output the assembled sequences after performing de novo assembly from raw signal overlaps ( 3). Rawsamble enables the use of existing de novo assemblers such as miniasm [68] as it provides the overlap information in a standardized format that these assemblers use.
@@ 2.2 Constructing an Index from Noisy Raw Signals via Aggressive Filtering
Rawsamble identifies overlapping regions between raw nanopore signals using a hash-based seeding mechanism that operates in two steps. First, to enable quick matching between raw signals, Rawsamble enables constructing an index directly from raw nanopore signals instead of using an existing reference genome sequence. While converting reference genomes to their expected raw signal values to store them in an index is mainly free from certain types of noise that raw nanopore signals contain (e.g., stochastic signal fluctuations [5] and variable speed of DNA molecules moving through nanopores [70, 71]), noise in raw nanopore signals can cause challenges to find accurate matches between raw signals when these signals are stored in an index. Second, to reduce noise stored in the hash tables and enable accurate similarity identification when both signals are noisy, Rawsamble aggressively filters raw signals as shown in Figure 2. The filtering mechanism iteratively compares two adjacent signals, $s_{i}$ and $s_{j}$, and removes the second signal, $s_{j}$, if the absolute difference between the two adjacent signals is below a certain threshold $T$. This filtering generates a list of filtered signals that aggressively aims to reduce the impact of stay errors during sequencing. Although similar filtering approaches are used in prior works [53, 58, 59] to reduce the stay errors (e.g., by aiming to perform homopolymer compression of raw signals [63]), Rawsamble employs a substantially larger threshold (i.e., aggressive) for filtering to minimize noise both in the index and during mapping. By applying the aggressive filtering technique, Rawsamble ensures that only high-quality, informative events are used in indexing to improve the accuracy and efficiency of the overall overlapping mechanism when both signals are noisy.
@@ 2.4 Adjusting the Mapping Strategy
Rawsamble adjusts the mapping strategy used in RawHash2 in two ways. First, Rawsamble generates pairwise mappings of a read to many reads (instead of a single mapping per read). To achieve this, Rawsamble identifies all valid chains based on their chaining scores and reports all such chains between raw signals. This enables overlapping a single raw signal with multiple other raw signals. Second, Rawsamble filters out trivial cyclic overlaps between two reads. To do so, Rawsamble avoids reporting overlapping signal pairs both as query and target in the mapping output, which can complicate the de novo assembly construction process [68, 74, 75]. To avoid these trivial cyclic overlaps, Rawsamble uses a pre-defined and deterministic ordering between raw signals (e.g., based on the lexicographic ordering of read names). By comparing raw signals deterministically, Rawsamble guarantees that only one of each pair of overlapping reads is processed as a query sequence while the other is treated as a target sequence. Reporting all chains while avoiding cyclic overlaps between raw signals allows for a comprehensive representation of the overlapping regions, which is useful for constructing accurate and long assemblies.
2.5 Signal Assembly from Overlaps
Rawsamble provides the overlapping information between raw signals using the Pairwise mApping Format (PAF) [68], which can be used by existing assemblers like miniasm [68] to construct an assembly as an assembly graph for output in Graphical Fragment Assembly (GFA) [68, 76] format.
To output the sequence of assembled signals directly from identified overlaps along with the assembly graph in GFA format, Rawsamble provides a new assembler, Rawasm [77]. This assembler can be useful for facilitating complete de novo assembly without requiring basecalling or to aid the basecalling process in several ways, as we discuss in Section 4.
Rawasm is implemented as a small extension to miniasm, introducing three key modifications to enable outputting the assembled signals. First, to read raw signals and store the assembled signals in standard raw signal file formats, Rawasm supports all major raw signal file formats: FAST5, POD5, S/BLOW5 [78]. Second, to improve memory efficiency and reduce the significant I/O burden associated with reading and writing raw signal files, which are typically larger and stored in complex compressed formats compared to basecalled sequences, Rawasm implements several optimizations. Unlike miniasm, which keeps all unitigs in main memory until the GFA file is written, Rawasm overlaps the graph building and writing processes. When the signals for a unitig are trimmed for outputting, and their positions within the unitig are calculated, Rawasm updates the output and frees up the corresponding memory. This approach minimizes memory usage and I/O delays to improve efficiency and reduce I/O overhead. Third, to output the assembled signals without losing the additional important metadata associated with raw signals, Rawasm generates a new raw signal file for each unitig in the assembly graph by concatenating the original signals that construct the corresponding unitig. The concatenation mechanism trims and arranges the signals based on the start and end positions used for constructing the unitig as provided by the assembly graph, similar to miniasm. Since each raw signal may contain different metadata (e.g., signal variables determined based on the characteristics of a nanopore channel), Rawasm keeps this metadata associated with each raw signal.
The overlapping information that Rawsamble generates can be used by any assembler that takes Pairwise mApping Format (PAF) [68] as input, including Rawasm and miniasm [68]. Although any such assembler can construct assemblies from PAF and output the resulting assembly information in GFA, Rawasm is the first implementation to enable outputting the resulting assembly signals directly after from the resulting GFA.
@@ 3.1 Evaluation Methodology
We implement the improvements we propose in Rawsamble on the RawHash2 implementation [59]. Similar to RawHash2 and minimap2 [73], Rawsamble provides the mapping information in the standard Pairwise mApping Format (PAF) [68]. We basecall the signals on two different hardware setups using Dorado’s 1) high accuracy (HAC) and super high accuracy (SUP) models with an NVIDIA RTX A6000 GPU [79], and 2) fast (Fast) and high accuracy (HAC) models with an Intel Xeon Gold 6226R CPU [80]. We use  Dorado’s 1) Fast model on a CPU to show the best performance (i.e., speed) that a CPU-based basecalling can provide and 2) SUP model on a GPU to show the best accuracy that basecalling provides along with its associated increased computational cost. We use POD5 files for all datasets, as suggested by Dorado [35] for optimal performance. Since no prior works can overlap reads using raw signals, we compare Rawsamble to minimap2, the current state-of-the-art read overlapper for basecalled sequences using the datasets shown in Table 1. We use datasets both with high sequencing depth of coverage (D1 to D3) and low coverage (D4 and D5) to evaluate the capability to construct assemblies in both scenarios.
@@
Organism Device Reads Bases Estimated SRA
Type (#) (#) Coverage ($\times$) Accession
D1 SARS-CoV-2 MinION 10,001 4.02M 135$\times$ CADDE Centre
D2 E. coli GridION 353,948 2,332M 445$\times$ ERR9127551
D3 Yeast MinION 50,023 385M 32$\times$ SRR8648503
D4 Green Algae PromethION 30,012 622M 5.6$\times$ ERR3237140
D5 Human MinION 270,006 1,773M 0.6$\times$ FAB42260
@@
All datasets are generated using ONT’s R9.4.1 e8 flow cells (450 bp/sec translocation speed with
4000 signals/sec sampling frequency).
All of the datasets are basecalled using Dorado (HAC) v3.3.
@@
We evaluate computational requirements in terms of overall run time with 32 CPU threads and peak memory usage. We use 32 CPU threads since the average thread utilization of CPU-based basecalling is around 32. We report the elapsed time, CPU time (when using only CPUs without GPUs), and peak memory usage of 1) minimap2 with and without basecalling and 2) Rawsamble. For Rawsamble, we additionally report throughput (average number of signals analyzed per second per single CPU thread) to provide insights about its capabilities for real-time analysis. To measure performance and peak memory usage, we use the time -v command in Linux when running Rawsamble, minimap2, and Dorado. For average speedup and memory comparisons of Rawsamble against other methods, we use the geometric mean to reduce the impact of outlier data points on the average calculation.
@@
We evaluate Rawsamble based on two use cases: 1) read overlapping and 2) de novo assembly. We generate read overlaps from 1) raw signals using Rawsamble and 2) basecalled sequences (with Dorado’s HAC model) of corresponding strands of raw signals using minimap2. To evaluate all-vs-all overlapping, we calculate the ratio of overlapping pairs 1) shared by both Rawsamble and minimap2, 2) unique to Rawsamble, and 3) unique to minimap2. For de novo assembly, we use miniasm [68] to assemble the read overlaps that Rawsamble and minimap2 generate in a PAF file. We use miniasm because it enables us to evaluate the impact of overlaps that Rawsamble and minimap2 finds by using the same assembler for both of them. To identify the roofline in terms of the assembly contiguity given a dataset, we use a highly accurate assembler as gold standard for our evaluations. To do this, we use Dorado’s most accurate model (i.e., SUP) to construct highly accurate reads and use Flye [81] to construct assemblies from these reads, as suggested in the guidelines by ONT [82]. This approach enables us to evaluate the gap between the assemblies that Rawsamble generates and the golden standard assembly that can be constructed from the same datasets.
@@
To evaluate the assemblies constructed from the overlaps that Rawsamble and minimap2 generate, we use several metrics that are mainly related to the contiguity of the assemblies. These metrics are 1) the total length of unitigs (Total Length), 2) the number of nucleotides in the largest assembly graph component (Largest Comp.), 3) unitig length at which 50% of the assembly is covered by unitigs of equal or greater length (N50), 4) area under the Nx curve to generate a more robust evaluation of contiguity (auN) [83], 5) longest unitig length, and 6) the number of unitigs. We use Bandage [84] to visualize these assemblies.
@@
We provide the parameter settings and versions for each tool as well as the details of the preset parameters in Rawsamble in Supplementary Tables S2 (parameters), S3 (details of presets), and S4 (versions). We provide the scripts to fully reproduce our results on the GitHub repository at https://github.com/CMU-SAFARI/RawHash, which contains the corresponding release version of Rawsamble we show in Supplementary Table S4. We provide the detailed reasoning behind the choice of Flye in order to generate the gold standard in Supplementary Material C.

### v2

@@ Abstract
Abstract: Raw nanopore signal analysis is a common approach in genomics to provide fast and resource-efficient analysis without translating the signals to bases (i.e., without basecalling). However, existing solutions cannot interpret raw signals directly if a reference genome is unknown due to a lack of accurate mechanisms to handle increased noise in pairwise raw signal comparison. Our goal is to enable the direct analysis of raw signals without a reference genome. To this end, we propose Rawsamble, the first mechanism that can identify regions of similarity between all raw signal pairs, known as all-vs-all overlapping, using a hash-based search mechanism.
We show that an off-the-shelf assembler (i.e., miniasm) can use the overlaps found by Rawsamble to construct genomes from scratch (i.e., de novo assembly). Our extensive evaluations across multiple genomes of varying sizes show that Rawsamble provides a significant speedup (on average by 4.48×4.48$\times$4.48 × and up to 23.10×23.10$\times$23.10 ×) and reduces peak memory usage (on average by 4.33×4.$3$ and up to by 22.00×22.$0$) compared to a conventional genome assembly pipeline using the state-of-the-art tools for basecalling (Dorado’s fastest mode) and overlapping (minimap2) on a CPU. We find that $30.94\%$ of overlapping pairs generated by Rawsamble are identical to those generated by minimap2. We further evaluate the benefits and directions that this new overlapping approach can enable in two ways. First, by constructing and evaluating the assembly graphs from the Rawsamble overlaps, we show that we can construct accurate and contiguous assembly segments (unitigs) up to 2.3 million bases in length (almost half the genome length of E. coli). Second, we identify previously unexplored directions that can be enabled by finding overlaps and constructing de novo assemblies as well as the challenges to tackle these future directions.
@@ 1 Introduction
Nanopore sequencing technology [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23] can sequence long nucleic acid molecules, known as reads, of up to a few million bases at high throughput [24, 25, 26]. As a molecule moves through a tiny pore, called a nanopore, ionic current measurements, called raw signals, are generated [5]. Nanopore sequencing provides three unique key benefits.
@@
Existing works that analyze raw nanopore signals [30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 27, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66] mainly utilize complex deep learning mechanisms [30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47] to translate these signals into nucleotides, a process called basecalling. Basecalling mechanisms usually 1) are designed to use large chunks of raw signal data for accurate analysis [53], 2) have high computational requirements [60, 26], and 3) oblivious to the similarity information between raw signals that lead to redundant basecalling on these similar regions. Existing basecalling techniques and the lack of understanding of similarity between raw signals can impose limitations to enable 1) accurate real-time analysis [53] and 2) portable sequencing with constrained resources [60] and prevent future work on designing a new class of basecalling techniques that can utilize the similarity information between raw signals.
To fully utilize the unique benefits of nanopore sequencing, it is necessary to analyze raw signals with 1) high accuracy and low latency for adaptive sampling and 2) low resource usage for portability and efficiency. To achieve this, several mechanisms focus on analyzing raw nanopore signals without basecalling [52, 27, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66] by mapping these signals to a reference genome. [^1] We use the raw signal analysis term specifically for these mechanisms in the remainder of the paper. Although raw nanopore signal mapping is widely studied [27, 53, 54, 57, 58, 59, 60, 62, 63, 64, 65, 66], none of these works without a reference genome to identify similarities directly between reads, called all-vs-all overlapping [67, 26, 68].
Although identifying all-vs-all overlapping between raw nanopore signals can be promising to enable new directions for both basecalling mechanisms (as we discuss in Section 4) and raw signal analysis (e.g., de novo assembly construction as we show in Section 3.4), it is challenging to do so for several reasons [5, 69, 70, 71]. First, similarities between a pair of signals must be identified accurately when both raw signals are noisy as compared to identifying similarities between a noisy raw signal and an accurate signal generated from a reference genome [53, 58, 59, 57]. Converting reference genomes to their expected raw signal values is free from certain types of noise that raw nanopore signals contain (e.g., stochastic signal fluctuations [5], variable speed of DNA molecules moving through nanopores [69, 70], and raw reads containing multiple split molecules), while raw signals are not free from such noise.
@@
Our goal is to enable 1) raw signal analysis without a reference genome and 2) new use cases with raw nanopore signal analysis by addressing the challenges of accurate and fast all-vs-all overlapping of raw nanopore signals. To this end, we propose Rawsamble, the first mechanism that enables fast and accurate overlap finding between raw nanopore signals. The key idea in Rawsamble is to re-design the existing state-of-the-art hash-based seeding mechanism [58, 59] for raw signals with more effective noise reduction techniques and useful outputting strategies to find all overlapping pairs accurately, which we explain in three key steps.
@@
  • We propose the first mechanism that can find all-vs-all overlapping of raw signals without basecalling to enable new use cases for raw signal analysis, such as assembly from overlapping raw signals and improving existing basecalling mechanisms.
  • We show that identifying overlaps with Rawsamble 1) is faster (on average 4.48×4.48$\times$4.48 × and up to 23.10×23.10$\times$23.10 ×) and 2) reduces peak memory requirements (on average 4.33×4.$3$ and up to 22.00×22.$0$) compared to the computational pipeline that includes the state-of-the-art basecalling running on a CPU with its fastest model followed by minimap2 to find overlaps. We evaluate the trade-offs between speed, accuracy, and hardware resources by running the basecaller on a CPU and GPU with fast and high-accuracy models.
  • We show that $30.94\%$ of overlapping pairs that Rawsamble generates are identical to the overlapping pairs generated by minimap2.
  • We report the first de novo assembly graphs ever constructed directly from raw signal overlaps without basecalling. We show that we can construct long unitigs up to 2.3 million bases in length for E. coli, which constitutes half the length of the genome. We find that these assemblies constructed from raw signal overlaps are either as accurate or more accurate than those created by minimap2 overlaps.
  • We identify new use cases that can be enabled by using overlapping raw signals and constructing assemblies from raw signals and discuss the advantages of generating them, such as for improving the accuracy and performance of basecalling. We discuss current limitations and challenges to enable these new use cases as future work.
@@ 2.1 Overview
Rawsamble is a mechanism to find overlapping pairs between raw nanopore signals (i.e., all-vs-all overlapping), which can be used by existing assemblers to construct de novo assembly graphs without basecalling, as shown in Figure 1. To achieve this, Rawsamble builds on the state-of-the-art raw signal mapper, RawHash2 [58, 59]. We provide the overview of RawHash2 in Supplementary Section A, where we explain the details related to processing raw signals such as generating events after the segmentation and generating hash values from these segmented raw signals (i.e., events). In the remainder of the paper, we use raw signals to refer to these events generated after segmentation.
Rawsamble extends RawHash2 to support all-vs-all raw signal overlapping in four key steps. First, to enable efficient and accurate indexing from the noisy raw signals ( 1), Rawsamble aggressively filters the raw input after the signal-to-event conversion to avoid nanopore-related errors. Second, to improve the accuracy of overlapping ( 2), Rawsamble adjusts the minimum chaining score to avoid false chains, ensuring that only high-confidence overlaps are considered. Third, to enable finding useful and long connections from many overlaps, Rawsamble adjusts the output strategy such that 1) all chains rather than only the best chain are reported and 2) cyclic overlaps are avoided. Fourth, Rawsamble enables the use of existing de novo assemblers off-the-shelf, such as miniasm [67], by providing the overlap information in a standardized format that these assemblers use.
@@ 2.2 Constructing an Index from Noisy Raw Signals via Aggressive Filtering
Rawsamble identifies overlapping regions between raw nanopore signals using a hash-based seeding mechanism that operates in two steps. First, to enable quick matching between raw signals, Rawsamble enables constructing an index directly from raw nanopore signals instead of using an existing reference genome sequence. While converting reference genomes to their expected raw signal values is mainly free from certain types of noise that raw nanopore signals contain (e.g., stochastic signal fluctuations [5] and variable speed of DNA molecules moving through nanopores [69, 70]), noise in raw nanopore signals can cause challenges to find accurate matches between raw signals when these signals are stored in an index. Second, to reduce noise stored in the hash tables and enable accurate similarity identification when both signals are noisy, Rawsamble aggressively filters raw signals as shown in Figure 2. The filtering mechanism iteratively compares two adjacent signals, $s_{i}$ and $s_{j}$, and removes the second signal, $s_{j}$, if the absolute difference between the two adjacent signals is below a certain threshold $T$. This filtering generates a list of filtered signals that aggressively aims to reduce the impact of stay errors during sequencing. Although similar filtering approaches are used in prior works [53, 58, 59] to reduce the stay errors (e.g., by aiming to perform homopolymer compression of raw signals [63]), Rawsamble employs a substantially larger threshold (i.e., aggressive) for filtering to minimize noise both in the index and during mapping. By applying the aggressive filtering technique, Rawsamble ensures that only high-quality, informative events are used in indexing to improve the accuracy and efficiency of the overall overlapping mechanism when both signals are noisy.
@@ 2.4 Adjusting the Mapping Strategy
Rawsamble adjusts the mapping strategy used in RawHash2 in two ways. First, Rawsamble generates pairwise mappings of a read to many reads (instead of a single mapping per read). To achieve this, Rawsamble identifies all valid chains based on their chaining scores and reports all such chains between raw signals. This enables the overlapping of a single raw signal with multiple other raw signals. Second, Rawsamble filters out trivial cyclic overlaps between two reads. To do so, Rawsamble avoids reporting overlapping signal pairs both as query and target in the mapping output, which can complicate the de novo assembly construction process [67, 73, 74]. To avoid these trivial cyclic overlaps, Rawsamble implements a pre-defined and deterministic ordering between raw signals (e.g., based on the lexicographic ordering of read names). By comparing raw signals deterministically, Rawsamble guarantees that only one of each pair of overlapping reads is processed as a query sequence while the other is treated as a target sequence. Reporting all chains while avoiding cyclic overlaps between raw signals allows for a comprehensive representation of the overlapping regions, which is useful for constructing accurate and long assemblies.
Output Format. Rawsamble provides the overlapping information between raw signals using the Pairwise mApping Format (PAF) [67]. To construct an assembly graph from the overlapping information that Rawsamble generates, any assembler that takes PAF files as input can be used, including miniasm [67].
@@ 3.1 Evaluation Methodology
We implement the improvements we propose in Rawsamble on the RawHash2 implementation [59]. Similar to RawHash2 and minimap2 [72], Rawsamble provides the mapping information in the standard Pairwise mApping Format (PAF) [67]. We basecall the signals on two different hardware setups using Dorado’s 1) high accuracy (HAC) model with an NVIDIA RTX A6000 GPU [75], and 2) fast (Fast) and HAC models with an Intel Xeon Gold 6226R CPU [76]. We include Dorado’s results on a CPU to provide a fair comparison on the same hardware platform, as providing the GPU implementation for Rawsamble is a future work. We use POD5 files for all datasets, as suggested by Dorado [35] for optimal performance. Since no prior works can overlap reads using raw signals, we compare Rawsamble to minimap2, the current state-of-the-art read overlapper for basecalled sequences using the datasets shown in Table 1. Although the R9.4.1 nanopore chemistry is deprecated, these raw datasets are commonly stored and basecalled for downstream analysis. We include datasets from both chemistries to show the feasibility and limitations of Rawsamble for each chemistry. For the R10.4.1 dataset, we use only the raw signals that contain a single molecule (i.e., non-chimeric multiple molecules can be contained in a single nanopore raw signal and later split by basecallers into multiple reads).
@@
Organism Chemistry Reads Bases Avg. Read Estimated SRA
Model (#) (#) Length Coverage ($\times$) or DOI
D1 E. coli R9.4 353,948 2,332M 6,588 445$\times$ ERR9127551
D2 Yeast R9.4 50,023 385M 7,698 32$\times$ SRR8648503
D3 Green Algae R9.4 30,012 622M TBC 5.6$\times$ ERR3237140
D4 E. coli R10.4.1 6,412,132 9,400M 1,466 1,797$\times$ 10.26188/25521892
@@
All of the datasets are basecalled using Dorado (HAC).
@@
We evaluate computational requirements in terms of overall run time with 64 CPU threads and peak memory usage. We use 64 CPU threads since the average thread utilization of CPU-based basecalling is around 64. We report the elapsed time, CPU time (when using only CPUs without GPUs), and peak memory usage of 1) Rawsamble and 2) basecalling followed by minimap2 (Dorado + Minimap2). For Rawsamble, we additionally report throughput (average number of signals analyzed per second per single CPU thread) to provide insights about its capabilities for real-time analysis. To measure performance and peak memory usage, we use the time -v command in Linux when running Rawsamble, minimap2, and Dorado. For average speedup and memory comparisons of Rawsamble against other methods, we use the geometric mean to reduce the impact of outlier data points on the average calculation.
@@
We evaluate Rawsamble based on two use cases: 1) read overlapping and 2) de novo assembly by constructing assembly graphs. We generate read overlaps from 1) raw signals using Rawsamble and 2) basecalled sequences (with Dorado’s HAC model) of corresponding strands of raw signals using minimap2. To evaluate the accuracy of all-vs-all overlapping, we calculate the ratio of overlapping pairs 1) shared by both Rawsamble and minimap2, 2) unique to Rawsamble, and 3) unique to minimap2. For de novo assembly, we use off-the-shelf miniasm [67] to generate assembly graphs from the read overlaps that Rawsamble and minimap2 generate. We use miniasm because it enables us to evaluate the impact of overlaps that Rawsamble and minimap2 find by using the same assembler for both of them. To identify the roofline in terms of the assembly contiguity given a dataset, we use a highly accurate assembler as gold standard for our evaluations. To do this, we use Dorado’s most accurate model (i.e., SUP) to construct highly accurate reads and use Flye [77] to construct assemblies from these reads, as suggested in the guidelines by ONT [78]. This approach enables us to evaluate the gap between the assemblies that Rawsamble generates and the golden standard assembly that can be constructed from the same datasets.
@@
To evaluate the contiguity of assemblies constructed from the overlaps that Rawsamble and minimap2 generate, we use several metrics. These metrics are 1) the total length of unitigs (Total Length), 2) the number of nucleotides in the largest assembly graph component (Largest Comp.), 3) unitig length at which 50% of the assembly is covered by unitigs of equal or greater length (N50), 4) area under the Nx curve to generate a more robust evaluation of contiguity (auN) [79], 5) longest unitig length, and 6) the number of unitigs. We estimate the basecalled read length of raw signals based on the sample frequency (i.e., the number of signals produced every second) and the translocation speed (i.e., the number of bases estimated to pass through a nanopore in a single second). Estimating the read length based on these parameters is a common and relatively consistent approach widely adopted in previous related works [54, 53, 58, 59]. We use Bandage [80] to visualize these assemblies. We note that our contiguity evaluations for Rawsamble and minimap2 are based on the assembly graphs constructed by miniasm.
[…]
@@
We provide the parameter settings and versions for each tool as well as the details of the preset parameters in Rawsamble in Supplementary Tables S2 (parameters), S3 (details of presets), and S4 (versions). We provide the scripts to fully reproduce our results on the GitHub repository at https://github.com/CMU-SAFARI/RawHash, which contains the corresponding release version of Rawsamble we show in Supplementary Table S4. We provide the detailed reasoning behind the choice of Flye in order to generate the gold standard in Supplementary Material B.